# Mitigating the Impact of Malware Evolution on API Sequence-based Windows Malware Detectors

### v1

@@ V-A Contrastive Learning Strategy
We design a contrastive learning strategy to enhances the encoder’s ability to capture fine-grained similarities and differences among API sequences and improve performance in detecting evolved malicious samples.
@@
The contrastive loss has two terms. The first term asks positive pairs from to be close together. These pairs are (benign, benign) or (malicious, malicious) pairs with the same malware family. In this way, the encoder will pay more attention to the similarities among samples in the same class. The evolved samples that retain API fragments similar to past samples will be represented as closer to past samples in the latent space. The second term aims to separate benign and malicious samples from each other, hopefully at least apart from each other. Thus, the encoder will focus on capturing the differences between benign and malicious samples, and prevent the classifier from misclassifying evolved malware as benign ones.
@@ VI-A Dataset
In this paper, we focus on malware of the Windows portable executable (PE) file which is the most popular malware file format. Our dataset, spanning over five years, contains 76,473 Windows PE files, i.e., 39,349 malicious and 37,124 benign as shown in Table IV. Specifically, The malicious software is obtained from the VirusShare website [31] and using a daily downloading script. The benign software is obtained from popular free software sources, including PortableApps [32], Softonic [33], SourceForge [34], and CNET [35].
To get reliable labels for these samples, we rely on VirusTotal[36] to determine whether a sample is benign or malicious. VirusTotal uses more than 60 anti-virus (AV) engines to vote whether the submitted sample is malicious or benign. In this paper, samples are labeled as malware when at least 10 AV engines report them as malicious, while samples are labeled as benign when no AV reports them as malicious. Note that according to a recent study [37] on measuring the labeling effectiveness of malware samples, this strategy is reasonable and stable. We consider samples up to Dec 2021 because following a previous work [38], the malware labels become stable after about one year, thus choosing Dec 2021 as the finishing time ensures good ground-truth confidence in objects labeled as malware.
Also, we leverage VirusTotal to get the exact appearing time for each sample and make sure that temporal consistency [13] is satisfied at the month level during the testing. Specifically, temporal consistency ensures that training samples should be strictly temporally precedent to testing ones, and all testing samples must come from the same period during each testing to eliminate time bias.
@@ VI-B API Sequence Extraction
After data collection, the Cuckoo Sandbox [39] is used to run the PE files and gather execution logs. Cuckoo sandbox has been widely used in prior works [7, 8, 9, 10]. It executes each PE file inside virtual machines and uses API hooks to monitor the Windows APIs to form a raw API sequence. In our system, dozens of virtual machines are maintained on the Cuckoo server which is installed with Ubuntu 16.04 LTS. All the virtual machines are installed with a 64-bit Windows 10 system and several necessary drivers to ensure the successful execution of the PE samples in the dataset. The snapshot feature of the virtual machine is leveraged to roll it back after execution to ensure the uniformity of the software running environment. Besides, Cuckoo simulates some user actions (such as clicking a button, typing some texts, etc.) to trigger malicious behavior of malware. In this paper, we set the maximum running time of each sample to 5 minutes. That is to say, the sandbox process completes when the uploaded sample ends itself or runs to 5 minutes. After a PE file is uploaded, Cuckoo server begins to call a free client to execute the file and record the API calls automatically. When the process completes, Cuckoo server will generate a sandbox report about this uploaded file and the raw API sequence can be extracted from this report.
@@ VI-C Evaluated Malware Detectors
We employ two representative DNNs, i.e., LSTM and Text-CNN, to build the malware detection models. These two models learn the sequence features from API sequences and have been proved to be effective in malware detection. In fact, many exsiting studies have already using these two models or their variants or combinations as the encoder [3, 4, 7, 5, 6]. The details of two DNN models are illustrated as follows:
@@ VI-C1 LSTM
LSTM [40] is a recurrent neural network architecture. It is able to capture the long-term context information through several gates designed to control the information transmission status. In this paper, we use the architecture of a single layer LSTM in [7] as our baseline model. Specifically, we establish two LSTM models for comparison, namely the regular LSTM model and the LSTM model enhanced by MME. The regular model includes an embedding layer [41] to receives API name sequences as input, a single layer LSTM encoder, and an MLP classifier. The enhanced LSTM includes our API sequence embedding, a contrastive LSTM encoder, and a classifier with the same configuration as the regular model.
@@ VI-C2 Text-CNN
Text-CNN [42] is a variant of CNN used for text classification tasks. The regular model here also use an embedding layer [41] and receives API name sequences as input. The filter size in CNN, or the n-gram size, denotes the number of successive API calls where the features are extracted. In the regular encoder, we set the filter sizes to 3, 4, and 5, respectively for three different Text-CNN layers. The enhanced Text-CNN includes our API sequence embedding, a contrastive Text-CNN encoder, and a classifier with the same configuration as the regular model.
@@ VII-A1 Experimental Settings
To evaluate the models’ sustainability, we test the mlaware detectors yearly. For each detectors, we train a model on the samples of 2017, and sequentially test its performance on each year from 2018 to 2021. To ensure the effectiveness of the models, we employ a 5-fold cross-validation during the model training process and ensure that the all the models achieve an average F1 score of over 97% on the validation set. During the model test, we calculate the false positive rate (FPR), false negitive rate (FNR), and F1 score to evaluate how MME can help prolong the life-time of regular models.
We also consider a state-of-the-art work called APIGraph [43], which is most relevant to our MME model, for comparison. APIGraph also leverages API knowledge graph learning and API clustering to enhance the regular malware detectors with capturing the semantically-equivalent APIs among evolved malware, thus slowing down the model aging. In fact, APIGraph primarily enhances the API name embedding stage of the model, whereas in comparison, our framework MME enhances both API name and argument embedding, as well as the encoder module.
@@ VII-A2 Results
Tabel V shows the performance of the each baseline model in every test year. The phenomenon of model aging is observed quite prominently, especially in terms of the FNR. Over a four-year testing period, the regular LSTM model exhibited an average FNR as high as 22.39%, resulting in a decrease in the F1 score to 83.98%. Similarly, the regular Text-CNN model showed an average FNR of 22.70%, with an accompanying drop in the F1 score to 84.67%. This indicates a severe issue of elevated false negatives caused by the evolution of malicious software, as the model tends to classify unknown malware as benign.
@@
| Testing Years | Regular LSTM | APIGraph(LSTM) | MME(LSTM) | ||||||
| FPR | FNR | F1 | FPR | FNR | F1 | FPR | FNR | F1 | |
| 2018 | 6.52 | 21.19 | 84.75 | 6.91 | 15.81 | 87.80 | 5.96 | 8.04 | 92.65 |
| 2019 | 6.59 | 17.59 | 86.25 | 7.18 | 15.79 | 86.96 | 7.27 | 7.37 | 91.61 |
| 2020 | 10.38 | 24.22 | 83.16 | 9.55 | 21.15 | 85.33 | 10.12 | 11.06 | 91.10 |
| 2021 | 8.96 | 26.57 | 81.78 | 8.15 | 23.24 | 84.19 | 8.62 | 14.12 | 89.55 |
| average | 8.11 | 22.39 | 83.98 | 7.94 | 18.99 | 86.07 | 7.99 | 10.15 | 91.23 |
| improve | – | – | – | 0.17 | 3.40 | 2.09 | 0.12 | 12.24 | 7.24 |
| Testing Years | Regular Text-CNN | APIGraph(Text-CNN) | MME(Text-CNN) | ||||||
| FPR | FNR | F1 | FPR | FNR | F1 | FPR | FNR | F1 | |
| 2018 | 5.19 | 19.45 | 86.50 | 5.62 | 12.90 | 90.13 | 2.39 | 6.72 | 95.23 |
| 2019 | 5.47 | 21.27 | 84.70 | 5.73 | 12.11 | 89.93 | 4.38 | 6.84 | 93.65 |
| 2020 | 5.87 | 24.97 | 83.98 | 7.18 | 17.01 | 88.54 | 5.39 | 10.86 | 92.62 |
| 2021 | 6.38 | 25.12 | 83.54 | 6.81 | 19.68 | 86.83 | 4.17 | 13.99 | 91.08 |
| average | 5.73 | 22.70 | 84.67 | 6.33 | 15.43 | 88.86 | 4.08 | 9.60 | 93.14 |
| improve | – | – | – | 0.61 | 7.27 | 4.18 | 1.65 | 13.10 | 8.47 |
@@ VII-B Model Maintainability Analysis
The purpose of this experiment is to evaluate how many human efforts MME can save while maintaining a high performance malware detection models.
Specifically, the comparison includes two aspects. On the one hand, we compare the amount of human efforts needed for active learning in maintaining both the regular and the enhanced models. On the other hand, we compare the model performance improvment given a fixed level of human effort.
@@ VII-B1 Comparison of human efforts needed to achieve a fixed performance
VII-B2 Comparison of model performance improvment given a fixed level of human efforts
The second comparison setting is to fix the amount of human efforts and test the model performance of the regular and enhanced models. Similarly, we train a detector with samples from 2012, and test the detector month by month from Jan 2018 to Dec 2021. We also use the uncertain sampling [13] in this experiment. We adopt two fixed human effort strategies: the first one is sample budgeting, where 20, 50, and 100 samples are labeled and used for retraining in each month; the second one is ratio budgeting, where 1%, 5%, and 10% of the samples from each month are labeled and used for retraining. Finally, we calculate the FPR, FNR, and F1 score of the model in each month, and calculate their respective averages as the final comparison metrics.
@@ VII-B2 Comparison of model performance improvment given a fixed level of human efforts
For both LSTM and Text-CNN, using the models enhanced with MME, only 20 samples or 1% of samples need to be labeled each month to keep the FNR below 10% and achieve an F1 score above 90%. In contrast, the regular models in our experiment require to label around 100 samples or 5% of the samples per month to achieve the same effect. This experimental result indicates that MME can reduce the analysts labeling effort by 5.
| Monthly Sample Budget | Base Model | Method | Average Performance | ||
| FPR(%) | FNR(%) | F1(%) | |||
| 20 | LSTM | Regular | 6.96 | 21.24 | 85.08 |
| MME | 7.73 | 9.19 | 91.86 | ||
| 0.77 | 12.05 | 6.77 | |||
| Text-CNN | Regular | 5.50 | 20.87 | 85.91 | |
| MME | 3.68 | 8.76 | 93.77 | ||
| 1.82 | 12.10 | 7.87 | |||
| 50 | LSTM | Regular | 5.42 | 15.68 | 89.05 |
| MME | 6.40 | 6.79 | 93.71 | ||
| 0.98 | 8.88 | 4.66 | |||
| Text-CNN | Regular | 2.54 | 15.99 | 90.10 | |
| MME | 3.28 | 7.93 | 94.40 | ||
| 0.74 | 8.06 | 4.30 | |||
| 100 | LSTM | Regular | 5.68 | 10.88 | 91.65 |
| MME | 4.80 | 5.83 | 94.90 | ||
| 0.88 | 5.05 | 3.25 | |||
| Text-CNN | Regular | 4.00 | 8.67 | 93.56 | |
| MME | 3.28 | 6.73 | 95.04 | ||
| 0.72 | 1.95 | 1.48 | |||
@@ VII-C1 Experimental Settings
In the MME framework we proposed, there are two enhanced components: embedding enhancement and encoder enhancement. The embedding enhancement consists of API name embedding in §III and API argument embedding in §IV. The encoder enhancement refers to the contrastive encoder in §V. To evaluate the impact of each component on the regular model’s enhancement, we construct two MME variants: one with only embedding enhancement and another with only encoder enhancement. We train the 4 models (one regular model, two MME variants, and one proposed MME enhanced model) on the samples of 2017, and test their performance on each year from 2018 to 2021. Based on the previous experiments, it is evident that the main indicators of decreased model performance are the increase in false negative rates and the decrease in F1 scores. Therefore, we use these two metrics to assess the influence of each part of the MME framework on the model’s enhancement. Figure 7 shows results of the ablation experiments, where the baseline models consist of LSTM and Text-CNN with the same experimental settings as §VI-C. The embedding enhanced LSTM/Text-CNN refers to the variant with only embedding enhancement, while the encoder enhanced LSTM/Text-CNN refers to the variant with only encoder enhancement.
| Monthly Ratio Budget | Base Model | Method | Average Performance | ||
| FPR(%) | FNR(%) | F1(%) | |||
| 1% | LSTM | Regular | 6.48 | 19.51 | 86.32 |
| MME | 6.66 | 8.35 | 92.76 | ||
| 0.18 | 11.16 | 6.44 | |||
| Text-CNN | Regular | 5.33 | 19.10 | 87.05 | |
| MME | 3.78 | 8.55 | 93.84 | ||
| 1.55 | 10.55 | 6.79 | |||
| 5% | LSTM | Regular | 6.41 | 9.92 | 91.86 |
| MME | 4.28 | 6.06 | 95.00 | ||
| 2.13 | 3.86 | 3.14 | |||
| Text-CNN | Regular | 3.97 | 8.46 | 93.69 | |
| MME | 3.64 | 6.03 | 95.25 | ||
| 0.34 | 2.43 | 1.56 | |||
| 10% | LSTM | Regular | 5.28 | 8.01 | 93.38 |
| MME | 3.90 | 5.45 | 95.49 | ||
| 1.39 | 2.56 | 2.11 | |||
| Text-CNN | Regular | 3.31 | 4.98 | 95.63 | |
| MME | 2.35 | 4.33 | 96.70 | ||
| 0.96 | 0.65 | 1.07 | |||

### v2

@@ V-A Contrastive Learning Strategy
We design a contrastive learning strategy to enhance the encoder’s ability to capture fine-grained similarities and differences among API sequences and improve performance in detecting evolved malicious samples.
@@
The contrastive loss has two terms. The first term asks positive pairs from to be close together. These pairs are (benign, benign) or (malicious, malicious) pairs with the same malware family. In this way, the encoder will pay more attention to the similarities among samples in the same class. The evolved samples that retain API fragments similar to past samples will be represented as closer to past samples in the latent space. The second term aims to separate benign and malicious samples from each other, hopefully at least apart from each other. Thus, the encoder will focus on capturing the differences between benign and malicious samples, and prevent the classifier from misclassifying evolved malware as benign samples.
@@ VI-A Dataset
In this paper, we focus on malware of the Windows portable executable (PE) file which is the most popular malware file format. Our dataset, spanning over five years, contains 76,473 Windows PE files, i.e., 39,349 malicious and 37,124 benign as shown in Table IV. Specifically, the malicious software is obtained from the VirusShare website [36] and using a daily downloading script. The benign software is obtained from popular free software sources, including PortableApps [37], Softonic [38], SourceForge [39], and CNET [40], BODMAS[41]. We made some of the data files publicly available111https://github.com/XingYuanWei/MME.
To get reliable labels for these samples, we rely on VirusTotal[42] to determine whether a sample is benign or malicious. VirusTotal uses more than 60 anti-virus (AV) engines to vote whether the submitted sample is malicious or benign. In this paper, samples are labeled as malware when at least 10 AV engines report them as malicious, while samples are labeled as benign when no AV engines report them as malicious. Note that according to a recent study [43] on measuring the labeling effectiveness of malware samples, this strategy is reasonable and stable. We consider samples up to Dec 2021 because following a previous work [44], the malware labels become stable after about one year, thus choosing Dec 2021 as the finishing time ensures good ground-truth confidence in objects labeled as malware.
Also, we leverage VirusTotal [42] to get the exact appearing time for each sample and make sure that temporal consistency [13] is satisfied at the month level during the testing. Specifically, temporal consistency ensures that training samples should be strictly temporally precedent to testing ones, and all testing samples must come from the same period during each testing to eliminate time bias.
@@ VI-B API Sequence Extraction
After data collection, the Cuckoo Sandbox [45] is used to run the PE files and gather execution logs. Cuckoo sandbox has been widely used in prior works [7, 8, 9, 10]. It executes each PE file inside virtual machines and uses API hooks to monitor the Windows APIs to form a raw API sequence. In our system, dozens of virtual machines are maintained on the Cuckoo server which is installed with Ubuntu 16.04 LTS. All the virtual machines are installed with a 64-bit Windows 10 system and several necessary drivers to ensure the successful execution of the PE samples in the dataset. The snapshot feature of the virtual machine is leveraged to roll it back after execution to ensure the uniformity of the software running environment. Besides, Cuckoo simulates some user actions (such as clicking a button, typing some text, etc.) to trigger the malicious behavior of malware. In this paper, we set the maximum running time of each sample to 5 minutes. That is to say, the sandbox process completes when the uploaded sample ends itself or runs for 5 minutes. After a PE file is uploaded, the Cuckoo server begins to call a free client to execute the file and record the API calls automatically. When the process completes, the Cuckoo server will generate a sandbox report about this uploaded file, and the raw API sequence can be extracted from this report.
@@ VI-C Evaluated Malware Detectors
We employ some representative DNNs, i.e., LSTM [26], Text-CNN [27], and Transformer [28] to build the malware detection models. These models learn the sequence features from API sequences and have been proven to be effective in malware detection. In fact, many existing studies have already used these three models or their variants or combinations as the encoder [3, 4, 7, 5, 6]. The details of two DNN models are illustrated as follows:
@@ VI-C1 LSTM
LSTM [26] is a recurrent neural network architecture. It is able to capture the long-term context information through several gates designed to control the information transmission status. In this paper, we use the architecture of a single-layer LSTM in [7] as our baseline model. Specifically, we establish two LSTM models for comparison, namely the regular LSTM model and the LSTM model enhanced by MME. The regular model includes an embedding layer [46] to receive API name sequences as input, a single layer LSTM encoder, and an MLP classifier. The enhanced LSTM includes our API sequence embedding, a contrastive LSTM encoder, and a classifier with the same configuration as the regular model.
Regular LSTM | APIGraph(LSTM)[21] | MME(LSTM) | ||||||||
| FPR | FNR | F1 | FPR | FNR | F1 | FPR | FNR | F1 | |||
| 2018 | 6.52 | 21.19 | 84.75 | 6.91 | 15.81 | 87.80 | 5.96 | 8.04 | 92.65 | ||
| 2019 | 6.59 | 17.59 | 86.25 | 7.18 | 15.79 | 86.96 | 7.27 | 7.37 | 91.61 | ||
| 2020 | 10.38 | 24.22 | 83.16 | 9.55 | 21.15 | 85.33 | 10.12 | 11.06 | 91.10 | ||
| 2021 | 8.96 | 26.57 | 81.78 | 8.15 | 23.24 | 84.19 | 8.62 | 14.12 | 89.55 | ||
| average | 8.11 | 22.39 | 83.98 | 7.94 | 18.99 | 86.07 | 7.99 | 10.15 | 91.23 | ||
| improve | – | – | – | 0.17 | 3.40 | 2.09 | 0.12 | 12.24 | 7.24 | ||
Regular Text-CNN | APIGraph(Text-CNN)[21] | MME(Text-CNN) | ||||||||
| FPR | FNR | F1 | FPR | FNR | F1 | FPR | FNR | F1 | |||
| 2018 | 5.19 | 19.45 | 86.50 | 5.62 | 12.90 | 90.13 | 2.39 | 6.72 | 95.23 | ||
| 2019 | 5.47 | 21.27 | 84.70 | 5.73 | 12.11 | 89.93 | 4.38 | 6.84 | 93.65 | ||
| 2020 | 5.87 | 24.97 | 83.98 | 7.18 | 17.01 | 88.54 | 5.39 | 10.86 | 92.62 | ||
| 2021 | 6.38 | 25.12 | 83.54 | 6.81 | 19.68 | 86.83 | 4.17 | 13.99 | 91.08 | ||
| average | 5.73 | 22.70 | 84.67 | 6.33 | 15.43 | 88.86 | 4.08 | 9.60 | 93.14 | ||
| improve | – | – | – | 0.61 | 7.27 | 4.18 | 1.65 | 13.10 | 8.47 | ||
Regular Transformer | APIGraph(Transformer)[21] | MME(Transformer) | ||||||||
| FPR | FNR | F1 | FPR | FNR | F1 | FPR | FNR | F1 | |||
| 2018 | 1.67 | 6.56 | 95.70 | 1.71 | 6.65 | 95.78 | 1.82 | 6.66 | 95.57 | ||
| 2019 | 3.13 | 7.23 | 94.23 | 3.13 | 7.37 | 95.01 | 2.63 | 6.91 | 94.72 | ||
| 2020 | 4.01 | 11.75 | 92.53 | 3.61 | 10.85 | 91.46 | 3.36 | 11.44 | 92.90 | ||
| 2021 | 3.06 | 15.51 | 90.56 | 2.74 | 15.81 | 89.48 | 2.73 | 14.81 | 91.08 | ||
| average | 2.97 | 10.26 | 93.25 | 2.80 | 10.17 | 92.93 | 2.63 | 9.95 | 93.57 | ||
| improve | – | – | – | 0.17 | 0.09 | 0.32 | 0.1 | 0.31 | 0.32 | ||
@@ VI-C2 Text-CNN
Text-CNN [27] is a variant of CNN used for text classification tasks. The regular model here also uses an embedding layer [46] and receives API name sequences as input. The filter size in CNN, or the n-gram size, denotes the number of successive API calls where the features are extracted. In the regular encoder, we set the filter sizes to 3, 4, and 5, respectively for three different Text-CNN layers. The enhanced Text-CNN includes our API sequence embedding, a contrastive Text-CNN encoder, and a classifier with the same configuration as the regular model.
VI-C3 Transformer
Transformer [28] is a neural network architecture based on self-attention mechanisms, originally designed for natural language processing tasks. It excels at capturing long-range dependencies and parallelizing computations, making it highly effective for sequence modeling. In this paper, we use a Transformer model as a baseline for comparison. The regular Transformer model includes an embedding layer that receives API name sequences as input, a Transformer encoder with multiple self-attention layers, and an MLP classifier. The enhanced Transformer model incorporates our API sequence embedding, a contrastive Transformer encoder, and a classifier with the same configuration as the regular model.
@@ VII-A1 Experimental Settings
To evaluate the models’ sustainability, we test the malware detectors yearly. For each detectors, we train a model on the samples of 2017, and sequentially test its performance on each year from 2018 to 2021. To ensure the effectiveness of the models, we employ a 10-fold cross-validation during the model training process and ensure that the all the models achieve an average F1 score of over 97% on the validation set. During the model test, we calculate the false positive rate (FPR), false negative rate (FNR), and F1 score to evaluate how MME can help prolong the life-time of regular models.
We also consider a state-of-the-art work called APIGraph [47],[21], which is most relevant to our MME model, for comparison. APIGraph also leverages API knowledge graph learning and API clustering to enhance the regular malware detectors by capturing the semantically-equivalent APIs among evolved malware, thus slowing down the model aging. In fact, APIGraph primarily enhances the API name embedding stage of the model, whereas in comparison, our framework MME enhances both API name and argument embedding, as well as the encoder module.
@@ VII-A2 Results
Table V shows the performance of each baseline model in every test year. The phenomenon of model aging is observed quite prominently, especially in terms of the FNR. Over a four-year testing period, the regular LSTM model exhibited an average FNR as high as 22.39%, resulting in a decrease in the F1 score to 83.98%. Similarly, the regular Text-CNN model showed an average FNR of 22.70%, with an accompanying drop in the F1 score to 84.67%. This indicates a severe issue of elevated false negatives caused by the evolution of malicious software, as the model tends to classify unknown malware as benign.
@@
The Transformer section of the table compares three models from 2018 to 2021. MME(Transformer) performs best, with the lowest average FPR (2.63%) and FNR (9.95%) and highest F1 (93.57%), improving FPR by 0.34%, FNR by 0.31%, and F1 by 0.32% over the regular Transformer (2.97% FPR, 10.26% FNR, 93.25% F1). APIGraph(Transformer) shows slight improvements, with 2.80% FPR, 10.17% FNR, and 92.93% F1, but is outperformed by MME(Transformer), which excels in reducing errors and boosting accuracy.
@@ VII-B Model Maintainability Analysis
The purpose of this experiment is to evaluate how many human efforts MME can save while maintaining a high-performance malware detection model.
Specifically, the comparison includes two aspects. On the one hand, we compare the amount of human effort needed for active learning in maintaining both the regular and the enhanced models. On the other hand, we compare the model performance improvement given a fixed level of human effort.
@@ VII-B1 Comparison of human efforts needed to achieve a fixed performance
VII-B2 Comparison of model performance improvement given a fixed level of human efforts
The second comparison setting is to fix the amount of human effort and test the model performance of the regular and enhanced models. Similarly, we train a detector with samples from 2012, and test the detector month by month from Jan 2018 to Dec 2021. We also use the uncertain sampling [13] in this experiment. We adopt two fixed human effort strategies: the first one is sample budgeting, where 20, 50, and 100 samples are labeled and used for retraining in each month; the second one is ratio budgeting, where 1%, 5%, and 10% of the samples from each month are labeled and used for retraining. Finally, we calculate the FPR, FNR, and F1 score of the model in each month, and calculate their respective averages as the final comparison metrics.
@@ VII-B2 Comparison of model performance improvement given a fixed level of human efforts
For both LSTM and Text-CNN, using the models enhanced with MME, only 20 samples or 1% of samples need to be labeled each month to keep the FNR below 10% and achieve an F1 score above 90%. In contrast, the regular models in our experiment require to label around 100 samples or 5% of the samples per month to achieve the same effect. This experimental result indicates that MME can reduce the analysts labeling effort by 5 times.
Method | Average Performance | ||||||
| FPR(%) | FNR(%) | F1(%) | |||||||
| 20 | LSTM | Regular | 6.96 | 21.24 | 85.08 | ||||
| MME | 7.73 | 9.19 | 91.86 | ||||||
| 0.77 | 12.05 | 6.77 | |||||||
| Text-CNN | Regular | 5.50 | 20.87 | 85.91 | |||||
| MME | 3.68 | 8.76 | 93.77 | ||||||
| 1.82 | 12.10 | 7.87 | |||||||
| 50 | LSTM | Regular | 5.42 | 15.68 | 89.05 | ||||
| MME | 6.40 | 6.79 | 93.71 | ||||||
| 0.98 | 8.88 | 4.66 | |||||||
| Text-CNN | Regular | 2.54 | 15.99 | 90.10 | |||||
| MME | 3.28 | 7.93 | 94.40 | ||||||
| 0.74 | 8.06 | 4.30 | |||||||
| 100 | LSTM | Regular | 5.68 | 10.88 | 91.65 | ||||
| MME | 4.80 | 5.83 | 94.90 | ||||||
| 0.88 | 5.05 | 3.25 | |||||||
| Text-CNN | Regular | 4.00 | 8.67 | 93.56 | |||||
| MME | 3.28 | 6.73 | 95.04 | ||||||
| 0.72 | 1.95 | 1.48 | |||||||
@@ VII-C1 Experimental Settings
In the MME framework we proposed, there are two enhanced components: embedding enhancement and encoder enhancement. The embedding enhancement consists of API name embedding in §III and API argument embedding in §IV. The encoder enhancement refers to the contrastive encoder in §V. To evaluate the impact of each component on the regular model’s enhancement, we construct two MME variants: one with only embedding enhancement and another with only encoder enhancement. We train the 4 models (one regular model, two MME variants, and one proposed MME enhanced model) on the samples of 2017, and test their performance on each year from 2018 to 2021. Based on the previous experiments, it is evident that the main indicators of decreased model performance are the increase in false negative rates and the decrease in F1 scores. Therefore, we use these two metrics to assess the influence of each part of the MME framework on the model’s enhancement. Figure 7 and Figure 8 show results of the ablation experiments, where the baseline models consist of LSTM, Text-CNN, and Transformer with the same experimental settings as §VI-C. The embedding enhanced LSTM, Text-CNN, and Transformer refer to the variant with only embedding enhancement, while the encoder enhanced LSTM/Text-CNN refers to the variant with only encoder enhancement.
Method | Average Performance | ||||||
| FPR(%) | FNR(%) | F1(%) | |||||||
| 1% | LSTM | Regular | 6.48 | 19.51 | 86.32 | ||||
| MME | 6.66 | 8.35 | 92.76 | ||||||
| 0.18 | 11.16 | 6.44 | |||||||
| Text-CNN | Regular | 5.33 | 19.10 | 87.05 | |||||
| MME | 3.78 | 8.55 | 93.84 | ||||||
| 1.55 | 10.55 | 6.79 | |||||||
| 5% | LSTM | Regular | 6.41 | 9.92 | 91.86 | ||||
| MME | 4.28 | 6.06 | 95.00 | ||||||
| 2.13 | 3.86 | 3.14 | |||||||
| Text-CNN | Regular | 3.97 | 8.46 | 93.69 | |||||
| MME | 3.64 | 6.03 | 95.25 | ||||||
| 0.34 | 2.43 | 1.56 | |||||||
| 10% | LSTM | Regular | 5.28 | 8.01 | 93.38 | ||||
| MME | 3.90 | 5.45 | 95.49 | ||||||
| 1.39 | 2.56 | 2.11 | |||||||
| Text-CNN | Regular | 3.31 | 4.98 | 95.63 | |||||
| MME | 2.35 | 4.33 | 96.70 | ||||||
| 0.96 | 0.65 | 1.07 | |||||||